OpenStudy (loveneenah):

Which statement about the order pairs 2,-9 and 3,-6 is true for this equation 5x-y/3=13? A.2,-9 B.3,-6 C. both are order pairs D.neither are orderd pair is a solution.

OpenStudy (loveneenah):

Could some one help

OpenStudy (goku_1999):

let's first ordered pair 2, -9. Now plug it in the equation: 5(2) - (-9/3) = 13 10 - (-3) = 13 10 + 3 = 13 13 = 13 Now the second ordered pair: 3, -6 5(3) - (-6/3) = 13 15 - (-2) = 13 15 + 2 = 13 17 = 13 false The ordered pair 2, -9 fits the equation so A

OpenStudy (loveneenah):

thank you!!

OpenStudy (goku_1999):

No problem, can I get a medal? :3

OpenStudy (loveneenah):

Yes you can get a medal!!

OpenStudy (loveneenah):

there u go!!

OpenStudy (loveneenah):

could u help me with a few more?

OpenStudy (loveneenah):

What is the x intercept of the line with this equation 8x-1/3y=16

OpenStudy (loveneenah):

Will fan + medal!

OpenStudy (goku_1999):

Just making sure is it -1 divided by 3y

OpenStudy (loveneenah):

So it is -0.3333ect.?

OpenStudy (goku_1999):

I got it, the x-intercept is 2. plot the order pair 2, 0 in the equation 8(2) - 1/3(0) = 16 16 - 0 = 16 16 = 16

OpenStudy (loveneenah):

thank you

OpenStudy (goku_1999):

No problem, need help with any more?

OpenStudy (loveneenah):

yes one more

OpenStudy (loveneenah):

ok?

OpenStudy (goku_1999):

Alright, what's the question?

OpenStudy (loveneenah):

what is y= -1/2x-4 on a graph

OpenStudy (goku_1999):

Do you need to describe what it looks like on a graph or do you need the image of the graph?

OpenStudy (loveneenah):

i need the image of the graph

OpenStudy (goku_1999):

Alright I'll work on the graph and send you a image

OpenStudy (loveneenah):

ok thanks!!

OpenStudy (goku_1999):

This it how it looks, if they are also looking for the x-intercept it is -8

OpenStudy (loveneenah):

Thank you

OpenStudy (loveneenah):

So it is -8 right?

OpenStudy (goku_1999):

You're welcome and it is -8 for the x intercept. I don't think they are asking for it so no need to worry about it

OpenStudy (loveneenah):

ok :)

OpenStudy (loveneenah):

It does ask for the y and the x intercept

OpenStudy (goku_1999):

The y intercept for the graph is -4. Ordered pair is (0, -4) The x intercept for the graph is -8. Ordered pair is (-8, 0)